What is the volume of 23.017g of butane gas (C4H10), at 26.7°C and 121.9 kPa? Round your answer to the nearest thousandth.  


1
Expert's answer
2021-07-09T05:15:29-0400

MM(C4H10) = 58.12 g/mol

n(C4H10) "= \frac{2}{MM}= \frac{23.017}{58.12}=0.396 \;mol"

Ideal gas law

"pV=nRT \\

V= \frac{nRT}{p} \\

p= 121.9 \;kPa \\

T= 26.7 + 273.15 = 299.85 \;K \\

R= 8.314 \;kPa \cdot L/K \cdot mol \\

V = \frac{0.396 \times 8.314 \times 299.85}{121.9}=8.098 \;L"

Answer: 8.098 L